OpenStudy (kyanda17):

Evaluate 3(a + b + c)2 squared for a = 2, b = 3, and c = 4.

OpenStudy (michele_laino):

hint: you can use this identity: \[\large {\left( {a + b + c} \right)^2} = {a^2} + {b^2} + {c^2} + 2ab + 2ac + 2bc\]

OpenStudy (michele_laino):

otherwise, you can substitute your numerical data into your expression: \[\Large \begin{gathered} a + b + c = 2 + 3 + 4 = 9 \hfill \\ 3{\left( {a + b + c} \right)^2} = 3 \cdot {9^2} = ...? \hfill \\ \end{gathered} \]